### v1

# Huan Zhang

Abstract
This paper develops a primal-dual dynamical system where the coefficients are designed in closed-loop way for solving a convex optimization problem with linear equality constraints. We first introduce a “second-order primal” + “first-order dual” continuous-time dynamical system, in which both the time scaling and Hessian-driven damping are governed by a feedback control of the gradient for the Lagrangian function. This system achieves the fast convergence rates for the primal-dual gap, the feasibility violation, and the objective residual along its trajectory. Subsequently, by time discretization of this system, we develop an accelerated primal-dual algorithm with a gradient-defined adaptive step size. We also obtain convergence rates for the primal-dual gap, the feasibility violation, and the objective residual. Furthermore, we provide numerical results to demonstrate the practical efficacy and superior performance of the proposed algorithm.
1 Introduction
1.1 Problem description
Let $\mathcal{H}$ and $\mathcal{G}$ be real Hilbert spaces equipped with inner product $\langle\cdot,\cdot\rangle$ and norm $\|\cdot\|$. In this paper, we consider the convex optimization problem with linear equality constraints of the form
$\left\{\begin{array}[]{ll}&\mathop{\mbox{min}}\limits_{x\in\mathcal{H}}~~{f(x)}\\ &\mbox{s.t.}~~Ax=b,\end{array}\right$. (3)
where $f:\mathcal{H}\rightarrow\mathbb{R}$ is a differentiable convex function with $L$-Lipschitz continuous gradient for $L>0$, $A:\mathcal{H}\rightarrow\mathcal{G}$ is a continuous linear operator and $b\in\mathcal{G}$. The optimization problem of the form (3) arises in many applications across fields such as image recovery, machine learning, and network optimization, see boyd2010 ; gold2014 ; ouyang2015 ; li2019 and the references therein.
The Lagrangian function associated with Problem (3) is defined as
$\mathcal{L}(x,\lambda):=f(x)+\left\langle\lambda,Ax-b\right\rangle,$
where $\lambda\in\mathcal{G}$ is the Lagrange multiplier. A pair $(x^{*},\lambda^{*})\in\mathcal{H}\times\mathcal{G}$ is said to be a saddle point of the Lagrangian function $\mathcal{L}$ iff
$\mathcal{L}(x^{*},\lambda)\leq\mathcal{L}(x^{*},\lambda^{*})\leq\mathcal{L}(x,\lambda^{*}),~~\forall(x,\lambda)\in\mathcal{H}\times\mathcal{G}.$
In the sequel, the set of saddle points of $\mathcal{L}$ is denoted by $\mathcal{S}$. The set of feasible points of Problem (3) is denoted by $\mathcal{F}:=\{x\in\mathcal{H}|Ax=b\}$. For any $(x,\lambda)\in\mathcal{F}\times\mathcal{G}$, it holds that $f(x)=\mathcal{L}(x,\lambda)$. We assume that $\mathcal{S}\neq\emptyset$. Let $(x^{*},\lambda^{*})\in\mathcal{S}$. Then,
$(x^{*},\lambda^{*})\in\mathcal{S}\Leftrightarrow\begin{cases}0=\nabla_{x}\mathcal{L}(x^{*},\lambda^{*})=\nabla f(x^{*})+A^{*}\lambda^{*},\\ 0=\nabla_{\lambda}\mathcal{L}(x^{*},\lambda^{*})=Ax^{*}-b,\end{cases}$
where $A^{*}:\mathcal{G}\rightarrow\mathcal{H}$ denotes the adjoint operator of $A$.
To solve Problem (3), in this paper, we propose the following “second-order primal” + “first-order dual” dynamical system whose damping is a feedback control of the gradient of the Lagrangian function:
$\begin{cases}\ddot{x}(t)+\frac{2[\dot{\tau}(t)]^{2}-\tau(t)\ddot{\tau}(t)}{\tau(t)\dot{\tau}(t)}\dot{x}(t)+\frac{[\dot{\tau}(t)]^{2}}{\tau(t)}\frac{d}{dt}\nabla_{x}\mathcal{L}\left(x(t),\lambda(t)\right)\\ ~~~~~~~~~~~~~~~~~~~~~~+\frac{\dot{\tau}(t)(\dot{\tau}(t)+\ddot{\tau}(t))}{\tau(t)}\nabla_{x}\mathcal{L}\left(x(t),\lambda(t)\right)=0,\\ \dot{\lambda}(t)-\dot{\tau}(t)\nabla_{\lambda}\mathcal{L}\left(x(t)+\frac{\tau(t)}{\dot{\tau}(t)}\dot{x}(t),\lambda(t)\right)=0,\\ \tau(t)-\frac{1}{q^{q}}\left(t_{0}+\int_{t_{0}}^{t}[\mu(s)]^{\frac{1}{q}}ds\right)^{q}=0,\\ [\mu(t)]^{p}\left\|\nabla_{x}\mathcal{L}\left(x(t),\lambda(t)\right)\right\|^{p-1}=1,\end{cases}$ (4)
where $q>0$, $p\geq 1$, $t\geq t_{0}>0$ and $\tau:[t_{0},+\infty)\rightarrow(0,+\infty)$ is the time scaling function which is non-decreasing and continuously differentiable. We obtain the fast convergence rates for the primal-dual gap, the feasibility violation, and the objective residual along the trajectory generated by System (4). Subsequently, we propose an autonomous primal-dual algorithm by time discretization of System (4), and give some convergence analysis.
1.2 Motivation and related works
Over the past decades, a series of studies have explored second-order dynamical systems in continuous and discrete time for the unconstrained convex optimization problem
$\mathop{\mbox{min}}\limits_{x\in\mathcal{H}}{f(x)}.$ (5)
Su et al. su2016 proposed the following second-order dynamical system:
$\ddot{x}(t)+\frac{\alpha}{t}\dot{x}(t)+\nabla f(x(t))=0,$ (6)
where $\alpha\geq 3$, and obtained the $\mathcal{O}\left(\frac{1}{t^{2}}\right)$ convergence rate of the objective residual along the trajectory generated by the system (6). The system (6) has further been studied in several papers, including may2017 ; attouch2018 ; attouch2019 . They showed that the objective residual converges at a rate of order $o\left(\frac{1}{t^{2}}\right)$ for $\alpha>3$, and the trajectory weakly converges to the global minimizer of the objective function. In order to improve the convergence rate, Attouch et al. a2019 proposed the system (6) with the time scaling:
$\ddot{x}(t)+\frac{\alpha}{t}\dot{x}(t)+\beta(t)\nabla f(x(t))=0,$
where $\alpha\geq 1$ and $\beta:[t_{0},+\infty)\rightarrow(0,+\infty)$ is the time scaling function which is non-decreasing and continuously differentiable. They obtained the $\mathcal{O}\left(\frac{1}{t^{2}\beta(t)}\right)$ convergence rate of the objective residual along the trajectory generated by this system, and proposed an inertial proximal algorithm by time discretization of this system.
On the other hand, primal-dual dynamics methods were also attracted an increasing interest by many researchers for solving the linear equality constrained convex optimization problem (3). Over the past few years, there have been a wide variety of works devoted to “second-order primal” + “second-order dual” continuous-time dynamical systems for solving Problem (3), see bot2021n ; hehu2021 ; zeng2023 ; hn2023 . By time discretization of continuous-time dynamical systems, new accelerated primal-dual algorithms for solving Problem (3) have been proposed in hehu2022 ; bot2023 ; zhu2024 ; dingliu2025 . Since the computational cost of a primal-dual method primarily arises from the sub-problem involving the primal variable, and a first-order dynamical system is simpler and more tractable to solve than a second-order one, He et al. hehufang2022 proposed the following “second-order primal” + “first-order dual” dynamical system for solving Problem (3):
$\begin{cases}\ddot{x}(t)+\gamma\dot{x}(t)+\beta(t)\nabla_{x}\mathcal{L}\left(x(t),\lambda(t)\right)=0,\\ \dot{\lambda}(t)-\beta(t)\nabla_{\lambda}\mathcal{L}\left(x(t)+\delta\dot{x}(t),\lambda(t)\right)=0,\end{cases}$
where $\gamma>0$ and $\delta>0$. This system can be viewed as an extension of Polyaks heavy ball with friction system in p1964 . Note that “second-order primal” + “first-order dual” dynamical systems and discrete accelerated algorithms for solving Problem (3) have been studied in 2024zhu ; li2025 ; zhufang2025 ; s2025 .
The present study is motivated by prior works in two domains: dynamical systems with Hessian-driven damping and those with closed-loop control. We now provide a comprehensive review of these research areas.
1.2.1 Dynamical systems with Hessian-driven damping
It is worth noting that the incorporation of Hessian-driven damping into the dynamical system has been a significant milestone in optimization and mechanics, as Hessian-driven damping can effectively mitigate the oscillations. Building on this foundation, many researchers use dynamical systems with Hessian-driven damping for solving the problem (5). Alvarez et al. a2002 first proposed the second-order dynamical system with constant viscous damping and constant Hessian-driven damping:
$\ddot{x}(t)+a\dot{x}(t)+\gamma\nabla^{2}f(x(t))\dot{x}(t)+b\nabla f(x(t))=0,$ (7)
where $b=1$. They obtained some new convergence properties of the solution trajectory generated by the system (7). Attouch et al. attouch2016 established the system (7) with $a=\frac{\alpha}{t}$, and obtained the $\mathcal{O}\left(\frac{1}{t^{2}}\right)$ convergence rate of the objective residual for solving the problem (5). To further accelerate the convergence rate, Attouch et al. a2022 proposed the system (7) with $a=\frac{\alpha}{t}$ and $b=\beta(t)$, and obtained the $\mathcal{O}\left(\frac{1}{t^{2}\beta(t)}\right)$ convergence rate for the objective residual. Moreover, for leading to Nesterov-type inertial algorithms via standard implicit or explicit schemes, Alecsa et al. Alecsa2021 proposed the second-order dynamical system with implicit Hessian-driven damping:
$\ddot{x}(t)+\frac{\alpha}{t}\dot{x}(t)+\nabla f\left(x(t)+\left(\gamma+\frac{\beta}{t}\right)\dot{x}(t)\right)=0.$
They showed that inertial algorithms, such as Nesterov’s algorithm, can be obtained via the natural explicit discretization from this dynamical system, and they also obtained the $\mathcal{O}\left(\frac{1}{t^{2}}\right)$ convergence rate of the objective residual along the trajectory generated by this system. For more results on second order dynamical systems with Hessian-driven damping for solving the problem (5), see bot2019 ; bot2021 ; attouch2023 and the references therein.
However, to the best of our knowledge, it appears that there exist only few papers in the literature devoting to the study of primal-dual dynamical systems with Hessian-driven damping in both continuous and discrete time for solving Problem (3). More precisely, He et al. hetian2025 proposed the following “second-order primal” + “first-order dual” dynamical system:
$\begin{cases}\ddot{x}(t)+\frac{\alpha}{t}\dot{x}(t)+\beta(t)\frac{d}{dt}\nabla_{x}\mathcal{L}\left(x(t),\lambda(t)\right)+\gamma(t)\nabla_{x}\mathcal{L}\left(x(t),\lambda(t)\right)=0,\\ \dot{\lambda}(t)-\eta(t)\nabla_{\lambda}\mathcal{L}\left(x(t)+\frac{t}{\alpha-1}\dot{x}(t),\lambda(t)\right)=0,\end{cases}$
where $t\geq t_{0}$, $\alpha>1$. They obtained the $\mathcal{O}\left(\frac{1}{t\eta(t)}\right)$ convergence rate for the primal-dual gap, feasibility violation and the objective residual along the trajectory generated by this system, and also obtained corresponding convergence rates for proposed primal-dual algorithm by time discretization of this system. Li et al. lihe2025 proposed a “second-order primal”+“second-order dual” dynamical system with implicit Hessian damping. They also established the fast convergence properties of the proposed dynamical system under suitable conditions. From what was mentioned above, we see that primal-dual dynamical systems with Hessian-driven damping for solving convex optimization problems with linear equality constraints have so far received much less attention compared with others. So, the first aim of this paper is to investigate primal-dual dynamical systems with Hessian-driven damping for solving Problem (3).
1.2.2 Dynamical systems with closed-loop control
The damping coefficients of a closed-loop dynamical system are directly governed by its evolving trajectory, which classifies the system as autonomous. Unlike non-autonomous systems, autonomous ones are generally preferred in practice because they do not depend on the time variable $t$. As far as we know, there are a few papers in the literature devoted to dynamical systems with closed-loop control for the unconstrained convex optimization problem (5). More precisely, Attouch et al. attouch2022 proposed an inertial dynamical system with closed-loop control:
$\ddot{x}(t)+\gamma(t)\dot{x}(t)+\nabla f(x(t))=0,$ (8)
where $t\geq t_{0}$ and $\gamma(t)=\|\dot{x}(t)\|^{p}$ for several values of $p\in\mathbb{R}$. They obtained the weak convergence of the trajectories to optimal solutions. For achieving convergence rate, Lin and Jordan lin2022 introduced a second-order dynamical system with closed-loop control as follows:
$\begin{cases}\ddot{x}(t)+\frac{2[\dot{\tau}(t)]^{2}-\tau(t)\ddot{\tau}(t)}{\tau(t)\dot{\tau}(t)}\dot{x}(t)+\frac{[\dot{\tau}(t)]^{2}}{\tau(t)}\nabla^{2}f(x(t))\dot{x}(t)+\frac{\dot{\tau}(t)(\dot{\tau}(t)+\ddot{\tau}(t))}{\tau(t)}\nabla f(x(t))=0,\\ \tau(t)-\frac{1}{4}\left(c+\int_{0}^{t}[\mu(s)]^{\frac{1}{2}}ds\right)^{2}=0,\\ [\mu(t)]^{p}\left\|\nabla f(x(t))\right\|^{p-1}=\theta,\end{cases}$ (9)
where $c>0$, $\theta\in(0,1)$, and $p\geq 1$. They obtained the $\mathcal{O}\left(\frac{1}{t^{\frac{3p+1}{2}}}\right)$ convergence rate for the objective residual along the trajectory generated by the system (9), and provided two algorithmic frameworks from time discretization of this system. Maier et al. m2023 proposed the system (8) with $\gamma(t):=\sqrt{E(t)}$, where $E(t):=f(x(t))-f(x^{*})+\frac{1}{2}\|\dot{x}(t)\|^{2}$, and obtained the $\mathcal{O}\left(\frac{1}{t^{2-\delta}}\right)$ convergence rate for the objective residual along the trajectory generated by this system with $\delta>0$. By using time scaling and averaging technique, Attouch et al. attouchbot2025 developed autonomous inertial dynamical systems which involve vanishing viscous damping and implicit Hessian-driven damping. They obtained the $\mathcal{O}\left(\frac{1}{t^{1+q-\frac{1}{p}}}\right)$ convergence rate for the objective residual along the trajectory generated by this system with $q>0$, $p\geq 1$ and the weak convergence of the trajectories to optimal solutions.
We observe that there is no results dealing with primal-dual dynamical systems with closed-loop control for solving Problem (3). So, the second aim of this paper is to investigate primal-dual dynamical systems with closed-loop control for solving Problem (3) in continuous and discrete time.
1.3 Main contributions
The contributions of this paper can be more specifically stated as follows:
The continuous level: We propose a new “second-order primal” + “first-order dual” dynamical system (4) with time scaling and Hessian-driven damping. The novelty of System (4) considered in this paper when compared with the system introduced in hehufang2022 is that System (4) incorporates Hessian-driven damping and all damping coefficients governed by the gradient of the Lagrangian function with a feedback control. We note that this is the first paper for the investigation of primal-dual dynamical systems with closed-loop control for solving convex optimization problems with linear equality constraints. Further, the convergence rate of the objective function residual obtained in this paper matches that reported in lin2022 , which addresses System (4) for unconstrained convex optimization problems.
The discrete level: We develop an accelerated autonomous primal-dual algorithm, where the step size is adaptively defined based on gradient of the objective function for Problem (3). This makes our numerical scheme much easier implementable than the numerical algorithm proposed in lin2022 . By appropriately adjusting these parameters, we show that this algorithm proposed in this paper exhibits the $\mathcal{O}\left(\frac{1}{k^{\frac{3p-1}{2p}}}\right)$ convergence rate for the primal-dual gap, the feasibility violation, and the objective residual. Moreover, in numerical experiments, our algorithm yields a significantly faster convergence rate and consistently achieves substantially higher accuracy than state-of-the-art methods.
1.4 Organization
The rest of this paper is organized as follows. In Section 2, we obtain fast convergence rates of the primal-dual gap, the feasibility violation, and the objective residual along the trajectories generated by System (4). In Section 3, we propose an accelerated autonomous primal-dual algorithm for solving Problem (3) and give some convergence analysis. In Section 4, we present some numerical experiments to illustrate the obtained results.
2 The continuous-time dynamical system
In this section, by using the Lyapunov analysis, we establish the fast convergence rates for the primal-dual gap, the feasibility violation, and the objective residual along the trajectory generated by System (4) under mild assumptions on the parameters. To begin, we recall the following result which will play an important role in the sequel.
Lemma 1.
(attouchbot2025, , Lemma 2.2) Suppose that there exist $C_{0}>0$ and $b>a\geq 0$ such that
$\int_{t_{0}}^{t}[\tau(s)]^{a}[\mu(s)]^{-b}ds\leq C_{0}<+\infty,~\forall t\geq t_{0},$
where $\tau$ is defined in terms of $\mu$ as in System (4), namely, $\tau(t)=\frac{1}{q^{q}}\left(t_{0}+\int_{t_{0}}^{t}[\mu(s)]^{\frac{1}{q}}ds\right)^{q}$. Then, there exists $C_{1}>0$ such that
$\tau(t)\geq C_{1}(t-t_{0})^{\frac{qb+1}{b-a}},~\forall t\geq t_{0}.$
Now, we investigate the convergence properties of System (4).
Theorem 2.1.
[…]
$\mathcal{L}(x(t),\lambda^{*})-\mathcal{L}(x^{*},\lambda^{*})=\mathcal{O}\left(\frac{1}{t^{\frac{2qp-p+1}{2}}}\right),~\textup{as}~t\to+\infty,$
$\|Ax(t)-b\|=\mathcal{O}\left(\frac{1}{t^{\frac{2qp-p+1}{2}}}\right),~\textup{as}~t\to+\infty,$
$|f(x(t))-f(x^{*})|=\mathcal{O}\left(\frac{1}{t^{\frac{2qp-p+1}{2}}}\right),~\textup{as}~t\to+\infty,$
and
$\int_{t_{0}}^{+\infty}[\dot{\tau}(s)]^{2}\left\|\nabla_{x}\mathcal{L}\left(x(s),\lambda(s)\right)\right\|^{2}ds<+\infty.$
Furthermore, $(x(s),\lambda(s))$ is bounded over $t_{0},t]$ and the upper bound only depends on the initial condition.
Proof.
Clearly, System (4) is equivalent to the following system:
$\begin{cases}\dot{v}(t)+\dot{\tau}(t)\nabla_{x}\mathcal{L}\left(x(t),\lambda(t)\right)=0,\\ \dot{x}(t)+\frac{\dot{\tau}(t)}{\tau(t)}(x(t)-v(t))+\frac{[\dot{\tau}(t)]^{2}}{\tau(t)}\nabla_{x}\mathcal{L}\left(x(t),\lambda(t)\right)=0,\\ \dot{\lambda}(t)-\dot{\tau}(t)\nabla_{\lambda}\mathcal{L}\left(x(t)+\frac{\tau(t)}{\dot{\tau}(t)}\dot{x}(t),\lambda(t)\right)=0,\\ \tau(t)-\frac{1}{q^{q}}\left(t_{0}+\int_{t_{0}}^{t}[\mu(s)]^{\frac{1}{q}}ds\right)^{q}=0,\\ [\mu(t)]^{p}\left\|\nabla_{x}\mathcal{L}\left(x(t),\lambda(t)\right)\right\|^{p-1}=1.\end{cases}$ (10)
For any fixed $(x^{*},\lambda^{*})\in\mathcal{S}$, we introduce the energy function $\mathcal{E}:\left[t_{0},+\infty\right)\to\mathbb{R}$ which defined as
$\begin{array}[]{rl}\mathcal{E}(t):=\tau(t)\left(\mathcal{L}(x(t),\lambda^{*})-\mathcal{L}(x^{*},\lambda^{*})\right)+\frac{1}{2}\lVert v(t)-x^{*}\rVert^{2}+\frac{1}{2}\lVert\lambda(t)-\lambda^{*}\rVert^{2}.\end{array}$ (11)
Obviously, $\mathcal{E}(t)\geq 0$ for all $t\geq t_{0}$. Note that
$\langle\dot{v}(t),v(t)-x^{*}\rangle$ $=\langle\dot{v}(t),v(t)-x(t)\rangle+\langle\dot{v}(t),x(t)-x^{*}\rangle$
$=\langle\dot{\tau}(t)\nabla_{x}\mathcal{L}\left(x(t),\lambda(t)\right),x(t)-v(t)\rangle-\langle\dot{\tau}(t)\nabla_{x}\mathcal{L}\left(x(t),\lambda(t)\right),x(t)-x^{*}\rangle.$
This together with (11) follows that
$\dot{\mathcal{E}}(t)=$ $\dot{\tau}(t)\left[\mathcal{L}\left(x(t),\lambda^{*}\right)-\mathcal{L}(x^{*},\lambda^{*})\right]+\tau(t)\langle\nabla_{x}\mathcal{L}(x(t),\lambda^{*}),\dot{x}(t)\rangle$
$+\langle\dot{v}(t),v(t)-x^{*}\rangle+\langle\dot{\lambda}(t),\lambda(t)-\lambda^{*}\rangle$
$=$ $\dot{\tau}(t)\left[\mathcal{L}\left(x(t),\lambda^{*}\right)-\mathcal{L}(x^{*},\lambda^{*})-\langle\nabla_{x}\mathcal{L}\left(x(t),\lambda(t)\right),x(t)-x^{*}\rangle\right]$
$+\tau(t)\langle\nabla_{x}\mathcal{L}(x(t),\lambda^{*}),\dot{x}(t)\rangle+\dot{\tau}(t)\langle\nabla_{x}\mathcal{L}\left(x(t),\lambda(t)\right),x(t)-v(t)\rangle$
$+\langle\dot{\lambda}(t),\lambda(t)-\lambda^{*}\rangle$
$=$ $\dot{\tau}(t)\left[\mathcal{L}\left(x(t),\lambda^{*}\right)-\mathcal{L}(x^{*},\lambda^{*})-\langle\nabla_{x}\mathcal{L}\left(x(t),\lambda^{*}\right),x(t)-x^{*}\rangle\right]$
$-\dot{\tau}(t)\langle\lambda(t)-\lambda^{*},Ax(t)-b\rangle+\tau(t)\langle\nabla_{x}\mathcal{L}(x(t),\lambda(t)),\dot{x}(t)\rangle$
$-\tau(t)\langle\lambda(t)-\lambda^{*},A\dot{x}(t)\rangle+\dot{\tau}(t)\langle\nabla_{x}\mathcal{L}\left(x(t),\lambda(t)\right),x(t)-v(t)\rangle$
$+\langle\dot{\lambda}(t),\lambda(t)-\lambda^{*}\rangle$
$=$ $\dot{\tau}(t)\left[\mathcal{L}\left(x(t),\lambda^{*}\right)-\mathcal{L}(x^{*},\lambda^{*})-\langle\nabla_{x}\mathcal{L}\left(x(t),\lambda^{*}\right),x(t)-x^{*}\rangle\right]$
$+\langle\nabla_{x}\mathcal{L}\left(x(t),\lambda(t)\right),\tau(t)\dot{x}(t)+\dot{\tau}(t)(x(t)-v(t))\rangle,$
where the third equality holds due to $\nabla_{x}\mathcal{L}\left(x(t),\lambda(t)\right)=\nabla_{x}\mathcal{L}\left(x(t),\lambda^{*}\right)+A^{*}(\lambda(t)-\lambda^{*})$ and the last equality holds due to the third equation of (10) and
$\nabla_{\lambda}\mathcal{L}\left(x(t)+\frac{\tau(t)}{\dot{\tau}(t)}\dot{x}(t),\lambda(t)\right)=A\left(x(t)+\frac{\tau(t)}{\dot{\tau}(t)}\dot{x}(t)\right)-b.$
By using the convexity of $\mathcal{L}(\cdot,\lambda^{*})$, we have
$\mathcal{L}\left(x(t),\lambda^{*}\right)-\mathcal{L}(x^{*},\lambda^{*})-\langle\nabla_{x}\mathcal{L}\left(x(t),\lambda^{*}\right),x(t)-x^{*}\rangle\leq 0.$
From the second equation of the system (10), we have
$\langle\nabla_{x}\mathcal{L}\left(x(t),\lambda(t)\right),\tau(t)\dot{x}(t)+\dot{\tau}(t)(x(t)-v(t))\rangle=-[\dot{\tau}(t)]^{2}\left\|\nabla_{x}\mathcal{L}\left(x(t),\lambda(t)\right)\right\|^{2}.$
Thus,
$\dot{\mathcal{E}}(t)\leq-[\dot{\tau}(t)]^{2}\left\|\nabla_{x}\mathcal{L}\left(x(t),\lambda(t)\right)\right\|^{2}\leq 0.$ (12)
This implies $\mathcal{E}(t)\leq\mathcal{E}(t_{0})$ for all $t\geq t_{0}$. By the definition of $\mathcal{E}(t)$, we obtain that
$\mathcal{L}\left(x(t),\lambda^{*}\right)-\mathcal{L}\left(x^{*},\lambda^{*}\right)\leq\frac{\mathcal{E}(t_{0})}{\tau(t)}=\mathcal{O}\left(\frac{1}{\tau(t)}\right),~\textup{as}~t\to+\infty,$
and $\lambda(t)$ is bounded on $t_{0},+\infty)$ and the upper bound only depends on the initial condition. From the third equation of the system (10), we have
$\lambda(t)-\lambda(t_{0})=\int_{t_{0}}^{t}\dot{\lambda}(s)ds$ $=\int_{t_{0}}^{t}\dot{\tau}(s)\left(Ax(s)-b+\frac{\tau(s)}{\dot{\tau}(s)}A\dot{x}(t)\right)ds$
$=\tau(t)(Ax(t)-b)-\tau(t_{0})(Ax(t_{0})-b).$
Then,
$\|\tau(t)(Ax(t)-b)\|<+\infty,$
which means that
$\|Ax(t)-b\|\leq\mathcal{O}\left(\frac{1}{\tau(t)}\right),~\textup{as}~t\to+\infty.$
From the definition of $\mathcal{L}$, we have
$|f(x(t))-f(x^{*})|\leq\mathcal{L}(x(t),\lambda^{*})-\mathcal{L}(x^{*},\lambda^{*})+\|\lambda^{*}\|\|Ax(t)-b\|\leq\mathcal{O}\left(\frac{1}{\tau(t)}\right),~\textup{as}~t\to+\infty.$
On the other hand, we deduce from (12) that
$\dot{\mathcal{E}}(t)+[\dot{\tau}(t)]^{2}\left\|\nabla_{x}\mathcal{L}\left(x(t),\lambda(t)\right)\right\|^{2}\leq 0.$
Then,
$\mathcal{E}(t)+\int_{t_{0}}^{t}[\dot{\tau}(s)]^{2}\left\|\nabla_{x}\mathcal{L}\left(x(s),\lambda(s)\right)\right\|^{2}ds\leq\mathcal{E}(t_{0}).$ (13)
It follows that
$\int_{t_{0}}^{+\infty}[\dot{\tau}(s)]^{2}\left\|\nabla_{x}\mathcal{L}\left(x(s),\lambda(s)\right)\right\|^{2}ds\leq\mathcal{E}(t_{0})<+\infty.$ (14)
From the fourth equation of the system (10), we have
$\dot{\tau}(t)=[\tau(t)]^{\frac{q-1}{q}}[\mu(t)]^{\frac{1}{q}}.$ (15)
This together with the last equation of the system (10) yields
$\dot{\tau}(t)]^{2}\left\|\nabla_{x}\mathcal{L}\left(x(t),\lambda(t)\right)\right\|^{2}=[\tau(t)]^{\frac{2q-2}{q}}[\mu(t)]^{\frac{2}{q}-\frac{2p}{p-1}}.$ (16)
Together with (14), (16) and Lemma 1, there exists $C_{1}>0$ such that
$\tau(t)\geq C_{1}(t-t_{0})^{\frac{2qp-p+1}{2}}.$ (17)
Thus,
$\mathcal{L}(x(t),\lambda^{*})-\mathcal{L}(x^{*},\lambda^{*})=\mathcal{O}\left(\frac{1}{t^{\frac{2qp-p+1}{2}}}\right),~\textup{as}~t\to+\infty,$
$\|Ax(t)-b\|=\mathcal{O}\left(\frac{1}{t^{\frac{2qp-p+1}{2}}}\right),~\textup{as}~t\to+\infty,$
and
$|f(x(t))-f(x^{*})|=\mathcal{O}\left(\frac{1}{t^{\frac{2qp-p+1}{2}}}\right),~\textup{as}~t\to+\infty.$
Now, we only need to show that $\tau(s)$ is bounded over $t_{0},t]$ and the upper bound only depends on the initial condition. In fact, by using the definition of $\mathcal{E}(t)$ and $\mathcal{E}(t)\leq\mathcal{E}(t_{0})$, we have
$\frac{1}{2}\|v(t)-x^{*}\|^{2}\leq\mathcal{E}(t_{0}).$
This follows that $v(t)$ is bounded over $t_{0},+\infty)$ and the upper bound only depends on the initial condition. Note that
$\tau(t)(x(t)-x^{*})-\tau(t_{0})(x(t_{0})-x^{*})=\int_{t_{0}}^{t}\left[\dot{\tau}(s)(x(s)-x^{*})+\tau(s)\dot{x}(s)\right]ds.$
Then,
$\|\tau(t)(x(t)-x^{*})\|\leq$ $\tau(t_{0})\|x(t_{0})-x^{*}\|+\int_{t_{0}}^{t}\left\|\dot{\tau}(s)(x(s)-x^{*})+\tau{(s)}\dot{x}(s)\right\|ds$ (18)
$\leq$ $\tau(t_{0})\|x(t_{0})-x^{*}\|+\int_{t_{0}}^{t}\|\dot{\tau}(s)(v(s)-x^{*})\|ds$
$+\int_{t_{0}}^{t}[\dot{\tau}(s)]^{2}\left\|\nabla_{x}\mathcal{L}\left(x(s),\lambda(s)\right)\right\|ds,$
where the last inequality holds due to the second equation of system (10). From (15) and the last equation of system (10), we have
$\dot{\tau}(t)]^{2}\left\|\nabla_{x}\mathcal{L}\left(x(t),\lambda(t)\right)\right\|=[\tau(t)]^{\frac{2q-2}{q}}\left\|\nabla_{x}\mathcal{L}\left(x(t),\lambda(t)\right)\right\|^{\frac{qp+2-2p}{qp}}.$ (19)
Combined with (18), (19), $\|v(t)-x^{*}\|\leq\sqrt{2\mathcal{E}(t_{0})}$ and the fact that $\tau(t)$ is monotonically increasing, we have
$\|x(t)-x^{*}\|\leq$ $\frac{\tau(t_{0})\|x(t_{0})-x^{*}\|+\sqrt{2\mathcal{E}(t_{0})}(\tau{(t)}-\tau(t_{0}))}{\tau{(t)}}$ (20)
$+\frac{1}{\tau{(t)}}\int_{t_{0}}^{t}\left\|[\tau(s)]^{\frac{2q-2}{q}}\nabla_{x}\mathcal{L}\left(x(s),\lambda(s)\right)^{\frac{qp+2-2p}{qp}}\right\|ds$
$\leq$ $\|x(t_{0})-x^{*}\|+\sqrt{2\mathcal{E}(t_{0})}$
$+\frac{1}{\tau(t)}\int_{t_{0}}^{t}[\tau(s)]^{\frac{2q-2}{q}}\left\|\nabla_{x}\mathcal{L}\left(x(s),\lambda(s)\right)\right\|^{\frac{qp+2-2p}{qp}}ds.$
Note that
$\frac{1}{\tau(t)}\int_{t_{0}}^{t}[\tau(s)]^{\frac{2q-2}{q}}\left\|\nabla_{x}\mathcal{L}\left(x(s),\lambda(s)\right)\right\|^{\frac{qp+2-2p}{qp}}ds$ (21)
$=$ $\frac{1}{\tau(t)}\int_{t_{0}}^{t}[\tau(s)]^{\frac{qp-2p}{2qp+2-2p}}[\tau(s)]^{\frac{qp}{2qp+2-2p}}\left([\tau(s)]^{\frac{2q-2}{q}}\left\|\nabla_{x}\mathcal{L}\left(x(s),\lambda(s)\right)\right\|^{\frac{2qp+2-2p}{qp}}\right)^{\frac{qp+2-2p}{2qp+2-2p}}ds.$
Now, we analyze (21) in terms of the following two cases.
$\mathbf{Case~I:}$ $1\leq q<2$ and $1\leq p\leq\frac{2}{2-q}$. In this case, from Hölder inequality and the fact that $\tau(t)$ is monotonically increasing, there exists $C_{2}>0$ such that
$\frac{1}{\tau(t)}\int_{t_{0}}^{t}[\tau(s)]^{\frac{qp-2p}{2qp+2-2p}}[\tau(s)]^{\frac{qp}{2qp+2-2p}}\left([\tau(s)]^{\frac{2q-2}{q}}\left\|\nabla_{x}\mathcal{L}\left(x(s),\lambda(s)\right)\right\|^{\frac{2qp+2-2p}{qp}}\right)^{\frac{qp+2-2p}{2qp+2-2p}}ds$ (22)
$\leq$ $[\tau(t_{0})]^{\frac{qp-2p}{2qp+2-2p}}\frac{1}{\tau(t)}\left(\int_{t_{0}}^{t}\tau(s)ds\right)^{\frac{qp}{2qp+2-2p}}\left(\int_{t_{0}}^{t}[\tau(s)]^{\frac{2q-2}{q}}\left\|\nabla_{x}\mathcal{L}\left(x(s),\lambda(s)\right)\right\|^{\frac{2qp+2-2p}{qp}}ds\right)^{\frac{qp+2-2p}{2qp+2-2p}}$
$\leq$ $C_{2}[\tau(t_{0})]^{\frac{qp-2p}{2qp+2-2p}}[\tau(t)]^{-\frac{qp+2-2p}{2qp+2-2p}}\left(\int_{t_{0}}^{t}[\tau(s)]^{\frac{2q-2}{q}}\left\|\nabla_{x}\mathcal{L}\left(x(s),\lambda(s)\right)\right\|^{\frac{2qp+2-2p}{qp}}ds\right)^{\frac{qp+2-2p}{2qp+2-2p}}$
$\leq$ $C_{2}[\tau(t_{0})]^{\frac{-2}{2qp+2-2p}}\left(\int_{t_{0}}^{t}[\tau(s)]^{\frac{2q-2}{q}}\left\|\nabla_{x}\mathcal{L}\left(x(s),\lambda(s)\right)\right\|^{\frac{2qp+2-2p}{qp}}ds\right)^{\frac{qp+2-2p}{2qp+2-2p}}.$
From (14) and (19), we have
$\small\int_{t_{0}}^{t}[\tau(s)]^{\frac{2q-2}{q}}\left\|\nabla_{x}\mathcal{L}\left(x(s),\lambda(s)\right)\right\|^{\frac{2qp+2-2p}{qp}}ds=\int_{t_{0}}^{t}[\dot{\tau}(s)]^{2}\left\|\nabla_{x}\mathcal{L}\left(x(s),\lambda(s)\right)\right\|^{2}ds\leq\mathcal{E}(t_{0}).$ (23)
Together with (20), (21), (22) and (23), we get
$\|x(t)-x^{*}\|\leq\|x(t_{0})-x^{*}\|+\sqrt{2\mathcal{E}(t_{0})}+C_{2}[\tau(t_{0})]^{\frac{-2}{2qp+2-2p}}[\mathcal{E}(t_{0})]^{\frac{qp+2-2p}{2qp+2-2p}}.$
$\mathbf{Case~II:}$ $q\geq 2$ and $p\geq 1$. In this case, from Hölder inequality and the fact that $\tau(t)$ is monotonically increasing, there exists $C_{3}>0$ such that
$\frac{1}{\tau(t)}\int_{t_{0}}^{t}[\tau(s)]^{\frac{qp-2p}{2qp+2-2p}}[\tau(s)]^{\frac{qp}{2qp+2-2p}}\left([\tau(s)]^{\frac{2q-2}{q}}\left\|\nabla_{x}\mathcal{L}\left(x(s),\lambda(s)\right)\right\|^{\frac{2qp+2-2p}{qp}}\right)^{\frac{qp+2-2p}{2qp+2-2p}}ds$
$\leq$ $[\tau(t)]^{\frac{-qp-2}{2qp+2-2p}}\left(\int_{t_{0}}^{t}\tau(s)ds\right)^{\frac{qp}{2qp+2-2p}}\left(\int_{t_{0}}^{t}[\tau(s)]^{\frac{2q-2}{q}}\left\|\nabla_{x}\mathcal{L}\left(x(s),\lambda(s)\right)\right\|^{\frac{2qp+2-2p}{qp}}ds\right)^{\frac{qp+2-2p}{2qp+2-2p}}$
$\leq$ $C_{3}[\tau(t)]^{\frac{-2}{2qp+2-2p}}\left(\int_{t_{0}}^{t}[\tau(s)]^{\frac{2q-2}{q}}\left\|\nabla_{x}\mathcal{L}\left(x(s),\lambda(s)\right)\right\|^{\frac{2qp+2-2p}{qp}}ds\right)^{\frac{qp+2-2p}{2qp+2-2p}}$
$\leq$ $C_{3}[\tau(t_{0})]^{\frac{-2}{2qp+2-2p}}\left(\int_{t_{0}}^{t}[\tau(s)]^{\frac{2q-2}{q}}\left\|\nabla_{x}\mathcal{L}\left(x(s),\lambda(s)\right)\right\|^{\frac{2qp+2-2p}{qp}}ds\right)^{\frac{qp+2-2p}{2qp+2-2p}}.$
This together with (20), (21) and (23) gives
$\|x(t)-x^{*}\|\leq\|x(t_{0})-x^{*}\|+\sqrt{2\mathcal{E}(t_{0})}+C_{3}[\tau(t_{0})]^{\frac{-2}{2qp+2-2p}}[\mathcal{E}(t_{0})]^{\frac{qp+2-2p}{2qp+2-2p}}.$
From both cases, it follows that $x(s)$ is bounded over $t_{0},t]$ and the upper bound only depends on the initial condition.
Remark 2.1.
Note that Attouch et al. attouchbot2025 introduced a second-order dynamical system via closed-loop control of the gradient for solving the problem (5). In (attouchbot2025, , Theorem 3.5), they obtained the ${o}\left(\frac{\ln(t)}{t^{pq}}\right)$ convergence rate for the objective residual along the trajectory generated by the system. Theorem 2.1 extends (attouchbot2025, , Theorem 3.5) from the unconstrained optimization problem to Problem (5), and achieves a faster convergence rate for the objective residual.
Now, let us consider special cases of System (4). In the special case when $q=1$, System (4) collapses to
$\begin{cases}\ddot{x}(t)+\frac{2[\dot{\tau}(t)]^{2}-\tau(t)\ddot{\tau}(t)}{\tau(t)\dot{\tau}(t)}\dot{x}(t)+\frac{[\dot{\tau}(t)]^{2}}{\tau(t)}\frac{d}{dt}\nabla_{x}\mathcal{L}\left(x(t),\lambda(t)\right)\\ ~~~~~~~~~~~~+\frac{\dot{\tau}(t)(\dot{\tau}(t)+\ddot{\tau}(t))}{\tau(t)}\nabla_{x}\mathcal{L}\left(x(t),\lambda(t)\right)=0,\\ \dot{\lambda}(t)-\dot{\tau}(t)\nabla_{\lambda}\mathcal{L}\left(x(t)+\frac{\tau(t)}{\dot{\tau}(t)}\dot{x}(t),\lambda(t)\right)=0,\\ \dot{\tau}(t)=\mu(t),\\ [\mu(t)]^{p}\left\|\nabla_{x}\mathcal{L}\left(x(t),\lambda(t)\right)\right\|^{p-1}=1.\end{cases}$ (24)
Corollary 1.
[…]
$\mathcal{L}(x(t),\lambda^{*})-\mathcal{L}(x^{*},\lambda^{*})=\mathcal{O}\left(\frac{1}{t^{\frac{p+1}{2}}}\right),~\textup{as}~t\to+\infty,$
$\|Ax(t)-b\|=\mathcal{O}\left(\frac{1}{t^{\frac{p+1}{2}}}\right),~\textup{as}~t\to+\infty,$
$|f(x(t))-f(x^{*})|=\mathcal{O}\left(\frac{1}{t^{\frac{p+1}{2}}}\right),~\textup{as}~t\to+\infty,$
and
$\int_{t_{0}}^{+\infty}[\dot{\tau}(s)]^{2}\left\|\nabla_{x}\mathcal{L}\left(x(s),\lambda(s)\right)\right\|^{2}ds<+\infty.$
Furthermore, $(x(s),\lambda(s))$ is bounded over $t_{0},t]$ and the upper bound only depends on the initial condition.
In the special case when $q=2$, System (4) collapses to the following system:
$\begin{cases}\ddot{x}(t)+\frac{2[\dot{\tau}(t)]^{2}-\tau(t)\ddot{\tau}(t)}{\tau(t)\dot{\tau}(t)}\dot{x}(t)+\frac{[\dot{\tau}(t)]^{2}}{\tau(t)}\frac{d}{dt}\nabla_{x}\mathcal{L}\left(x(t),\lambda(t)\right)\\ ~~~~~~~~~~~~~+\frac{\dot{\tau}(t)(\dot{\tau}(t)+\ddot{\tau}(t))}{\tau(t)}\nabla_{x}\mathcal{L}\left(x(t),\lambda(t)\right)=0,\\ \dot{\lambda}(t)-\dot{\tau}(t)\nabla_{\lambda}\mathcal{L}\left(x(t)+\frac{\tau(t)}{\dot{\tau}(t)}\dot{x}(t),\lambda(t)\right)=0,\\ \tau(t)-\frac{1}{4}\left(t_{0}+\int_{t_{0}}^{t}[\mu(s)]^{\frac{1}{2}}ds\right)^{2}=0,\\ [\mu(t)]^{p}\left\|\nabla_{x}\mathcal{L}\left(x(t),\lambda(t)\right)\right\|^{p-1}=1.\end{cases}$ (25)
Corollary 2.
[…]
$\mathcal{L}(x(t),\lambda^{*})-\mathcal{L}(x^{*},\lambda^{*})=\mathcal{O}\left(\frac{1}{t^{\frac{3p+1}{2}}}\right),~\textup{as}~t\to+\infty,$
$\|Ax(t)-b\|=\mathcal{O}\left(\frac{1}{t^{\frac{3p+1}{2}}}\right),~\textup{as}~t\to+\infty,$
$|f(x(t))-f(x^{*})|=\mathcal{O}\left(\frac{1}{t^{\frac{3p+1}{2}}}\right),~\textup{as}~t\to+\infty,$
and
$\int_{t_{0}}^{+\infty}[\dot{\tau}(s)]^{2}\left\|\nabla_{x}\mathcal{L}\left(x(s),\lambda(s)\right)\right\|^{2}ds<+\infty.$
Furthermore, $(x(s),\lambda(s))$ is bounded over $t_{0},t]$ and the upper bound only depends on the initial condition.
Remark 2.2.
Note that Lin and Michael lin2022 introduced the second-order dynamical system with closed-loop control (9). In (lin2022, , Theorem 3), they obtained
$f(x(t))-f(x^{*})=\mathcal{O}\left(\frac{1}{t^{\frac{3p+1}{2}}}\right),~\textup{as}~t\to+\infty.$
Therefore, Corollary 2 extends (lin2022, , Theorem 3) from the unconstrained optimization problem to Problem (5).
In the special case for System (4) when $p=q=1$. This implies that $\mu(t)=1$ and $\tau(t)=t$. System (4) collapses to the following open-loop system:
$\begin{cases}\ddot{x}(t)+\frac{2}{t}\dot{x}(t)+\frac{1}{t}\frac{d}{dt}\nabla_{x}\mathcal{L}\left(x(t),\lambda(t)\right)+\frac{1}{t}\nabla_{x}\mathcal{L}\left(x(t),\lambda(t)\right)=0,\\ \dot{\lambda}(t)-\nabla_{\lambda}\mathcal{L}\left(x(t)+t\dot{x}(t),\lambda(t)\right)=0.\end{cases}$ (26)
Corollary 3.
[…]
$\mathcal{L}(x(t),\lambda^{*})-\mathcal{L}(x^{*},\lambda^{*})=\mathcal{O}\left(\frac{1}{t}\right),~\textup{as}~t\to+\infty,$
$\|Ax(t)-b\|=\mathcal{O}\left(\frac{1}{t}\right),~\textup{as}~t\to+\infty,$
$|f(x(t))-f(x^{*})|=\mathcal{O}\left(\frac{1}{t}\right),~\textup{as}~t\to+\infty,$
and
$\int_{t_{0}}^{+\infty}\left\|\nabla_{x}\mathcal{L}\left(x(s),\lambda(s)\right)\right\|^{2}ds<+\infty.$
Furthermore, $(x(s),\lambda(s))$ is bounded over $t_{0},t]$ and the upper bound only depends on the initial condition.
Remark 2.3.
The system (26) can be as a special case of the system proposed in hetian2025 . Clearly, the parameters of the system (26) satisfy the Assumption 2.1 in hetian2025 . Thus, by Theorem 2.3 in hetian2025 , we have
$\mathcal{L}(x(t),\lambda^{*})-\mathcal{L}(x^{*},\lambda^{*})=\mathcal{O}\left(\frac{1}{t}\right),~\textup{as}~t\to+\infty,$
$\|Ax(t)-b\|=\mathcal{O}\left(\frac{1}{t}\right),~\textup{as}~t\to+\infty,$
and
$|f(x(t))-f(x^{*})|=\mathcal{O}\left(\frac{1}{t}\right),~\textup{as}~t\to+\infty,$
which is compatible with the results in Corollary 3.
3 The discrete accelerated autonomous primal-dual algorithm
In this section, we propose an accelerated autonomous primal-dual algorithm with adaptive step size defined via gradient and analyze the convergence properties of this algorithm. For simplicity, we restrict our discussion to the case where $q=1$, which gives $\dot{\tau}(t)=\mu(t)$. Obviously, the system (24) can be reformulated as:
$\begin{cases}y(t)=x(t)+\frac{\tau(t)}{\dot{\tau}(t)}\dot{x}(t),\\ \dot{y}(t)=-\dot{\tau}(t)\frac{d}{dt}\nabla_{x}\mathcal{L}\left(x(t),\lambda(t)\right)-(\dot{\tau}(t)+\ddot{\tau}(t))\nabla_{x}\mathcal{L}\left(x(t),\lambda(t)\right),\\ \dot{\lambda}(t)=\dot{\tau}(t)\left(Ay(t)-b\right),\\ \dot{\tau}(t)=\mu(t),\\ [\mu(t)]^{p}\left\|\nabla_{x}\mathcal{L}\left(x(t),\lambda(t)\right)\right\|^{p-1}=1.\end{cases}$ (27)
For the system (27), we consider the time step size $1$ and set $y(k)\approx y_{k}$, $x(k)\approx x_{k}$, $\tau(k)\approx\tau_{k}$, $\dot{\tau}(k)\approx\gamma_{k}$, $\lambda(k)\approx\lambda_{k}$, and $\mu(k)\approx\mu_{k}$. Evaluating the terms $\frac{d}{dt}\nabla_{x}\mathcal{L}\left(x(t),\lambda(t)\right)$ and $\ddot{\tau}(t)$ as the time derivative of $\nabla_{x}\mathcal{L}\left(x(t),\lambda(t)\right)$ and $\dot{\tau}(t)$, respectively. Now, by the time discretization of the system (27), we have
$\begin{cases}y_{k}=x_{k}+\frac{\tau_{k}}{\gamma_{k}}(x_{k}-x_{k-1}),\\ y_{k+1}-y_{k}=-\gamma_{k}\left[\nabla f(x_{k+1})+A^{*}\lambda_{k+1}-(\nabla f(x_{k})+A^{*}\lambda_{k})\right]\\ \qquad\qquad\qquad-(\gamma_{k+1}+\gamma_{k+1}-\gamma_{k})\left(\nabla f(x_{k+1})+A^{*}\lambda_{k+1}\right),\\ \lambda_{k+1}-\lambda_{k}=\gamma_{k+1}\left(Ay_{k+1}-b\right),\\ \tau_{k+1}=\gamma_{k}+\tau_{k},\\ \gamma_{k+1}=\mu_{k},\\ \mu_{k}^{p}\left\|\nabla_{x}\mathcal{L}\left(x_{k},\lambda_{k}\right)\right\|^{p-1}=1.\end{cases}$ (28)
Substituting the first equation and the third equation into the second equation in (28), we have
$\nabla f(x_{k+1})+\frac{\gamma_{k+1}+\tau_{k+1}}{2\gamma_{k+1}^{2}}(x_{k+1}-\bar{x}_{k})+(\gamma_{k+1}+\tau_{k+1})A^{*}(Ax_{k+1}-\sigma_{k+1})=0,$
where
$\bar{x}_{k}:=x_{k}+\frac{\gamma_{k+1}}{\gamma_{k+1}+\tau_{k+1}}\left(\frac{\tau_{k}}{\gamma_{k}}(x_{k}-x_{k-1})+\gamma_{k}\left(\nabla f(x_{k})+A^{*}\lambda_{k}\right)\right),$
and
$\sigma_{k+1}:=\frac{1}{\gamma_{k+1}+\tau_{k+1}}\left(\tau_{k+1}Ax_{k}+\gamma_{k+1}b-\lambda_{k}\right).$
This implies that
$x_{k+1}=$ $\mathop{\arg\min}\limits_{x\in\mathcal{H}}\left\{f(x)+\frac{\gamma_{k+1}+\tau_{k+1}}{4\gamma_{k+1}^{2}}\|x-\bar{x}_{k}\|^{2}+\frac{\gamma_{k+1}+\tau_{k+1}}{2}\left\|Ax-\sigma_{k+1}\right\|^{2}\right\}.$
Based on the above analysis, we are now in the position to introduce the following accelerated autonomous primal-dual algorithm for solving Problem (3).
Algorithm 1 Accelerated autonomous primal-dual algorithm (AAPDA)
Initialization: Choose $x_{0}=x_{1}\in\mathcal{H}\textup{ and }\lambda_{1}\in\mathcal{G}$. Let $p>1$, $\tau_{1}=0$ and $\gamma_{1}\geq 1$.
for $k=1,2,\dots$ do
  if
$\nabla_{x}\mathcal{L}\left(x_{k},\lambda_{k}\right)=0,$
then
   stop
  else
   Step 1: Compute
$\mu_{k}=\left\|\nabla_{x}\mathcal{L}\left(x_{k},\lambda_{k}\right)\right\|^{-\frac{p-1}{p}},$ (29)
$\gamma_{k+1}=\mu_{k},$ (30)
$\tau_{k+1}=\gamma_{k}+\tau_{k},$ (31)
$\bar{x}_{k}:=x_{k}+\frac{\gamma_{k+1}}{\gamma_{k+1}+\tau_{k+1}}\left(\frac{\tau_{k}}{\gamma_{k}}(x_{k}-x_{k-1})+\gamma_{k}\left(\nabla f(x_{k})+A^{*}\lambda_{k}\right)\right),$ (32)
$\sigma_{k+1}:=\frac{1}{\gamma_{k+1}+\tau_{k+1}}\left(\tau_{k+1}Ax_{k}+\gamma_{k+1}b-\lambda_{k}\right).$ (33)
   Step 2: Update the primal variable
$x_{k+1}=$ $\mathop{\arg\min}\limits_{x\in\mathcal{H}}\left\{f(x)+\frac{\gamma_{k+1}+\tau_{k+1}}{4\gamma_{k+1}^{2}}\|x-\bar{x}_{k}\|^{2}+\frac{\gamma_{k+1}+\tau_{k+1}}{2}\left\|Ax-\sigma_{k+1}\right\|^{2}\right\}.$ (34)
   Step 3: Compute
$y_{k+1}=x_{k+1}+\frac{\tau_{k+1}}{\gamma_{k+1}}(x_{k+1}-x_{k}),$ (35)
and update the dual variable
$\lambda_{k+1}=\lambda_{k}+\gamma_{k+1}(Ay_{k+1}-b).$ (36)
  end if
end for
In order to analyze the convergence rates of AAPDA, we need the following result.
Lemma 2.
(attouchbot2025, , Lemma 4.2) Let $\{\mu_{k}\}_{k\geq 0}$ be a positive sequence and $\{\tau_{k}\}_{k\geq 1}$ such that $\tau_{1}=0$ and $\tau_{k}=\sum_{i=0}^{k-2}\mu_{i}$ for all $k\geq 2$. Suppose that there exist $C_{4}>0$ and $a,b,c\geq 0$ such that $b+c>a$ and
$\sum_{k\geq 0}\tau_{k}^{a}\mu_{k}^{-b}\mu_{k+1}^{-c}\leq C_{4}<+\infty.$
Then, there exists $C_{5}>0$ such that for every $k\geq 2$, it holds $\tau_{k+1}\geq C_{5}k^{\frac{b+c+1}{b+c-a}}.$
The following equality will be used in the sequel:
$\frac{1}{2}\|x_{1}\|^{2}-\frac{1}{2}\|x_{2}\|^{2}=\langle x_{1},x_{1}-x_{2}\rangle-\frac{1}{2}\|x_{1}-x_{2}\|^{2},~~\forall x_{1},x_{2}\in\mathcal{H}.$ (37)
Now, we prove the convergence rates of the primal-dual gap, the feasibility violation, and the objective residual generated by AAPDA.
Theorem 3.1.
Let $\{(x_{k},\lambda_{k})\}_{k\geq 0}$ be the sequence generated by $\rm{AAPDA}$. Suppose that $\{\mu_{k}\}_{k\geq 0}$ is a non-decreasing sequence satisfying $\mu_{k}\geq 1$ for all $k\geq 0$. Also, let $\{\tau_{k}\}_{k\geq 1}$ be a sequence such that $\tau_{1}=0$ and $\tau_{k}=\sum_{i=0}^{k-2}\mu_{i}$ for all $k\geq 2$. Then, for any $\left(x^{*},\lambda^{*}\right)\in\mathcal{S}$, we have
$\mathcal{L}(x_{k},\lambda^{*})-\mathcal{L}(x^{*},\lambda^{*})=\mathcal{O}\left(\frac{1}{k^{\frac{3p-1}{2p}}}\right),\textup{ as }k\rightarrow+\infty,$
$\|Ax_{k}-b\|=\mathcal{O}\left(\frac{1}{k^{\frac{3p-1}{2p}}}\right),\textup{ as }k\rightarrow+\infty,$
$|f(x_{k})-f(x^{*})|=\mathcal{O}\left(\frac{1}{k^{\frac{3p-1}{2p}}}\right),\textup{ as }k\rightarrow+\infty,$
and
$\sum_{k=0}^{+\infty}\gamma_{k+1}^{2}\lVert\nabla_{x}\mathcal{L}(x_{k+1},\lambda_{k+1})\rVert^{2}<+\infty.$
Proof.
For $\left(x^{*},\lambda^{*}\right)\in\mathcal{S}$, we introduce the energy sequence which defined as
$E_{k}:=\tau_{k+1}\left(\mathcal{L}(x_{k},\lambda^{*})-\mathcal{L}(x^{*},\lambda^{*})\right)+\frac{1}{2}\lVert u_{k}\rVert^{2}+\frac{1}{2}\lVert\lambda_{k}-\lambda^{*}\rVert^{2},$
where $u_{k}:=y_{k}-x^{*}+\gamma_{k}\nabla_{x}\mathcal{L}(x_{k},\lambda_{k})$.
Clearly, from (32), (33), (34), (35) and (36) (see also the second equation of system (28)), it is easy to show that
$y_{k+1}-y_{k}=\gamma_{k}\nabla_{x}\mathcal{L}(x_{k},\lambda_{k})-2\gamma_{k+1}\nabla_{x}\mathcal{L}(x_{k+1},\lambda_{k+1}).$
Then,
$u_{k+1}-u_{k}$ $=y_{k+1}-y_{k}+\gamma_{k+1}\nabla_{x}\mathcal{L}(x_{k+1},\lambda_{k+1})-\gamma_{k}\nabla_{x}\mathcal{L}(x_{k},\lambda_{k})$
$=-\gamma_{k+1}\nabla_{x}\mathcal{L}(x_{k+1},\lambda_{k+1}).$
This together with (37) and $\nabla_{x}\mathcal{L}\left(x_{k+1},\lambda_{k+1}\right)=\nabla_{x}\mathcal{L}\left(x_{k+1},\lambda^{*}\right)+A^{*}(\lambda_{k+1}-\lambda^{*})$ gives
$\frac{1}{2}\lVert u_{k+1}\rVert^{2}-\frac{1}{2}\lVert u_{k}\rVert^{2}$ (38)
$=$ $\langle u_{k+1}-u_{k},u_{k+1}\rangle-\frac{1}{2}\lVert u_{k+1}-u_{k}\rVert^{2}$
$\leq$ $-\gamma_{k+1}\langle\nabla_{x}\mathcal{L}(x_{k+1},\lambda_{k+1}),y_{k+1}-x^{*}\rangle-\gamma_{k+1}^{2}\lVert\nabla_{x}\mathcal{L}(x_{k+1},\lambda_{k+1})\rVert^{2}$
$=$ $-\gamma_{k+1}\langle\nabla_{x}\mathcal{L}(x_{k+1},\lambda^{*}),y_{k+1}-x^{*}\rangle-\gamma_{k+1}\langle A^{*}(\lambda_{k+1}-\lambda^{*}),y_{k+1}-x^{*}\rangle$
$-\gamma_{k+1}^{2}\lVert\nabla_{x}\mathcal{L}(x_{k+1},\lambda_{k+1})\rVert^{2}.$
From (35) and the convexity of $\mathcal{L}(\cdot,\lambda^{*})$, we have
$-\gamma_{k+1}\langle\nabla_{x}\mathcal{L}(x_{k+1},\lambda^{*}),y_{k+1}-x^{*}\rangle$ (39)
$=$ $-\gamma_{k+1}\langle\nabla_{x}\mathcal{L}(x_{k+1},\lambda^{*}),x_{k+1}-x^{*}\rangle-\tau_{k+1}\langle\nabla_{x}\mathcal{L}(x_{k+1},\lambda^{*}),x_{k+1}-x_{k}\rangle$
$\leq$ $-\gamma_{k+1}(\mathcal{L}(x_{k+1},\lambda^{*})-\mathcal{L}(x^{*},\lambda^{*}))-\tau_{k+1}(\mathcal{L}(x_{k+1},\lambda^{*})-\mathcal{L}(x_{k},\lambda^{*}))$
$=$ $-(\gamma_{k+1}+\tau_{k+1})(\mathcal{L}(x_{k+1},\lambda^{*})-\mathcal{L}(x^{*},\lambda^{*}))+\tau_{k+1}(\mathcal{L}(x_{k},\lambda^{*})-\mathcal{L}(x^{*},\lambda^{*})).$
On the other hand, it follows from (36) and (37) that
$-\gamma_{k+1}\langle A^{*}(\lambda_{k+1}-\lambda^{*}),y_{k+1}-x^{*}\rangle$ $=-\gamma_{k+1}\langle\lambda_{k+1}-\lambda^{*},Ay_{k+1}-b\rangle$ (40)
$=-\langle\lambda_{k+1}-\lambda^{*},\lambda_{k+1}-\lambda_{k}\rangle$
$=-\frac{1}{2}\left(\lVert\lambda_{k+1}-\lambda^{*}\rVert^{2}-\lVert\lambda_{k}-\lambda^{*}\rVert^{2}+\lVert\lambda_{k+1}-\lambda_{k}\rVert^{2}\right).$
Together with (38)–(40), we have
$E_{k+1}-E_{k}$ $=\tau_{k+2}\left(\mathcal{L}(x_{k+1},\lambda^{*})-\mathcal{L}(x^{*},\lambda^{*})\right)-\tau_{k+1}\left(\mathcal{L}(x_{k},\lambda^{*})-\mathcal{L}(x^{*},\lambda^{*})\right)$ (41)
$~~~~+\frac{1}{2}\lVert u_{k+1}\rVert^{2}-\frac{1}{2}\lVert u_{k}\rVert^{2}+\frac{1}{2}\lVert\lambda_{k+1}-\lambda^{*}\rVert^{2}-\frac{1}{2}\lVert\lambda_{k}-\lambda^{*}\rVert^{2}$
$\leq(\tau_{k+2}-\tau_{k+1}-\gamma_{k+1})\left(\mathcal{L}(x_{k+1},\lambda^{*})-\mathcal{L}(x^{*},\lambda^{*})\right)$
$~~~~-\gamma_{k+1}^{2}\lVert\nabla_{x}\mathcal{L}(x_{k+1},\lambda_{k+1})\rVert^{2}-\frac{1}{2}\lVert\lambda_{k+1}-\lambda_{k}\rVert^{2}$
$\leq 0,$
where the last inequality holds due to (31). Thus, the sequence $\{E_{k}\}_{k\geq 0}$ is non-increasing. This implies that
$\mathcal{L}(x_{k},\lambda^{*})-\mathcal{L}(x^{*},\lambda^{*})=\mathcal{O}\left(\frac{1}{\tau_{k+1}}\right),\textup{ as }k\rightarrow+\infty,$ (42)
and $\{\lambda_{k}\}_{k\geq 0}$ is bounded.
For notation simplicity, denote a sequence $\{g_{k}\}_{k\geq 0}$ by
$g_{k}:=\tau_{k+1}(Ax_{k}-b).$
It follows from (31), (35) and (36) that
$\lambda_{k+1}-\lambda_{0}$ $=\sum_{j=0}^{k}(\lambda_{j+1}-\lambda_{j})$
$=\sum_{j=0}^{k}\gamma_{j+1}(Ay_{j+1}-b)$
$=\sum_{j=0}^{k}\left[(\gamma_{j+1}+\tau_{j+1})(Ax_{j+1}-b)-\tau_{j+1}(Ax_{j}-b)\right]$
$=\sum_{j=0}^{k}(g_{j+1}-g_{j})$
$=g_{k+1}-g_{0}.$
This together with the boundedness of $\{\lambda_{k}\}_{k\geq 0}$ follows that $\sup_{k\geq 0}\|g_{k}\|<+\infty$. Then,
$\|Ax_{k}-b\|\leq\mathcal{O}\left(\frac{1}{\tau_{k+1}}\right),\textup{ as }k\rightarrow+\infty.$ (43)
Combining (42), (43) and the definition of $\mathcal{L}$, we have
$\left|f(x_{k})-f(x^{*})\right|\leq\mathcal{L}(x_{k},\lambda^{*})-\mathcal{L}(x^{*},\lambda^{*})+\|\lambda^{*}\|\|Ax_{k}-b\|\leq\mathcal{O}\left(\frac{1}{\tau_{k+1}}\right),\textup{ as }k\rightarrow+\infty.$ (44)
Summing (41) over $k=0,\dots,N$, we obtain
$\sum_{k=0}^{N}\gamma_{k+1}^{2}\lVert\nabla_{x}\mathcal{L}(x_{k+1},\lambda_{k+1})\rVert^{2}\leq E_{0}-E_{N}\leq E_{0}.$
Letting $N\rightarrow+\infty$ in the above inequality, we obtain
$\sum_{k=0}^{+\infty}\gamma_{k+1}^{2}\lVert\nabla_{x}\mathcal{L}(x_{k+1},\lambda_{k+1})\rVert^{2}<+\infty.$
This together with (29) and (30) gives
$\sum_{k=0}^{+\infty}\mu_{k}^{2}\mu_{k+1}^{-\frac{2p}{p-1}}<+\infty.$
From $\mu_{k}\geq 1$ for all $k\geq 0$, we have
$\sum_{k=0}^{+\infty}\mu_{k+1}^{-\frac{2p}{p-1}}\leq\sum_{k=0}^{+\infty}\mu_{k}^{2}\mu_{k+1}^{-\frac{2p}{p-1}}<+\infty.$
Then, by using Lemma 2, there exists $C_{3}>0$ such that
$\tau_{k+1}\geq C_{3}k^{\frac{3p-1}{2p}}.$
This together with (42), (43) and (44) follows that
$\mathcal{L}(x_{k},\lambda^{*})-\mathcal{L}(x^{*},\lambda^{*})=\mathcal{O}\left(\frac{1}{k^{\frac{3p-1}{2p}}}\right),\textup{ as }k\rightarrow+\infty,$
$\|Ax_{k}-b\|=\mathcal{O}\left(\frac{1}{k^{\frac{3p-1}{2p}}}\right),\textup{ as }k\rightarrow+\infty,$
and
$|f(x_{k})-f(x^{*})|=\mathcal{O}\left(\frac{1}{k^{\frac{3p-1}{2p}}}\right),\textup{ as }k\rightarrow+\infty.$
The proof is complete.
4 Numerical experiments
In this section, we validate the theoretical findings of previous sections through numerical experiments. In these experiments, all codes are implemented in MATLAB R2021b and executed on a PC equipped with a 2.40GHz Intel Core i5-1135G7 processor and 16GB of RAM.
Example 4.1.
Consider the following constrained minimization problem
$\left\{\begin{array}[]{ll}&\mathop{\mbox{min}}\limits_{x\in\mathbb{R}^{n}}\frac{\mu}{2}\|x\|^{2}\\ &\mbox{s.t.}~~Ax=b,\end{array}\right$.
where $\mu\geq 0$, $A\in\mathbb{R}^{m\times n}$ and $b\in\mathbb{R}^{m}$. We generate the matrix $A$ using the standard Gaussian distribution. The original solution $x^{*}\in\mathbb{R}^{n}$ is generated from a Gaussian distribution $\mathcal{N}(0,4)$, with its entries clipped to the interval [-2,2] and sparsified so that only $1\%$ of its elements are non-zero. Furthermore, we select $b=Ax^{*}$.
(a) $n=m=10$
(b) $n=m=300$
(c) $n=m=2000$
Figure 1: Numerical results of AAPDA, FPDA and AALM under different dimensions.
Let $\mu=1.5$. The stopping condition is
$\frac{\left\|x_{k+1}-x_{k}\right\|}{\max\{\left\|x_{k}\right\|,1\}}\leq 10^{-6}$
or the number of iterations exceeds 100.
In the following experiments, we compare AAPDA with the fast primal-dual algorithm (FPDA) proposed in hehu2022 and the accelerated linearized augmented Lagrangian method (AALM) proposed in xu2017 . Here are parameter settings of algorithms:
  • AAPDA: $\gamma_{1}=1$, $\tau_{1}=0$ and $p=\{4,5,k\}$.
  • FPDA: $\alpha=50$, $\theta=2$, $\beta_{0}=\frac{0.2}{\theta}$ and $\epsilon_{k}=0$.
  • AALM: $\gamma=0.1$, $\alpha_{k}=\frac{2}{k+1}$, $\beta_{k}=\gamma_{k}=k\gamma$ and $P_{k}=\frac{1}{k}\mathrm{Id}$.
As shown in Figure 1, AAPDA achieves significantly superior performance over FPDA and AALM under all three different dimensions. It is worth noting that the convergence speed and accuracy of AAPDA improve with an increase in the value of $p$, and AAPDA can achieve exponential convergence. Moreover, compared to FPDA, Hessian-driven damping term in AAPDA can induces significant attenuation of the oscillations.
Example 4.2.
Consider the following non-negative least squares problem
$\mathop{\mbox{min}}\limits_{x\in\mathbb{R}^{n}}f(x):=\frac{1}{2}\|Ax-b\|^{2},$
where $A\in\mathbb{R}^{m\times n}$ and $b\in\mathbb{R}^{m}$. We generate a random matrix $A\in\mathbb{R}^{m\times n}$ with density $s\in(0,1]$ and a random $b\in\mathbb{R}^{m}$. The nonzero entries of $A$ are independently generated from a uniform distribution in $[0,0.1]$. The original solution $x^{*}$ is obtained via $x^{*}:=A\setminus b$ in MATLAB.
Let $m=500$ and $n=1000$. The stopping condition is
$\frac{\left\|x_{k+1}-x_{k}\right\|}{\max\{\left\|x_{k}\right\|,1\}}\leq\theta$
or the number of iterations exceeds 200. For $s=0.5$ and $s=1$, we respectively consider $\theta:=\{10^{-6},10^{-8},10^{-10}\}$.
In the following experiments, we compare AAPDA with FISTA proposed in beck2009 , the proximal inertial algorithm (PIA) proposed in (attouchbot2025, , Algorithm 4) and the accelerated forward-backward method (AFBM) proposed in a2016 . Here are the parameter settings of algorithms:
  • AAPDA: $\gamma_{1}=5$, $\tau_{1}=0$ and $p=5$.
  • FISTA: $\alpha=\frac{1}{\|A\|^{2}}$ and $t_{1}=1$.
  • PIA: $p=5$ and $\tau_{1}=0$.
  • AFBM: $\alpha=\frac{1}{\|A\|^{2}}$, $t_{1}=1$ and $\beta=5$.
(a) $s=0.5$
(b) $s=1$
Figure 2: Numerical results of AAPDA, FISTA, PIA and AFBM when $\theta=10^{-6}$.
(a) $s=0.5$
(b) $s=1$
Figure 3: Numerical results of AAPDA, FISTA, PIA and AFBM when $\theta=10^{-8}$.
(a) $s=0.5$
(b) $s=1$
Figure 4: Numerical results of AAPDA, FISTA, PIA and AFBM when $\theta=10^{-10}$.
As shown in Figures 2, 3 and 4, AAPDA demonstrates superior convergence performance compared to FISTA, PIA and AFBM. Specifically, AAPDA achieves a significantly faster convergence rate and consistently reaches a much higher final accuracy across different values of $\theta$. Moreover, AAPDA exhibits more stable performance, while the convergence speed and final accuracy of FISTA, PIA and AFBM are obviously affected by $\theta$.
5 Conclusion
In this paper, for solving a convex optimization problem with linear equality constraints (3), we propose a primal-dual dynamical system (4), and obtain the fast convergence rates for the primal-dual gap, the feasibility violation, and the objective residual along the trajectory of System (4). Specifically, the damping in System (4) is a feedback control of the gradient of the Lagrangian function for Problem (3). We also propose an accelerated autonomous primal-dual algorithm (AAPDA), which is derived from the time discretization of System (4), and establish convergence rates for the primal-dual gap, the feasibility violation, and the objective residual.
As shown in this paper, the convergence of the trajectories generated by the primal-dual dynamical system with closed-loop control is still a challenge. To address this, future work will reset the primal-dual dynamical system with closed-loop control and the energy function to achieve the convergence of the trajectories generated by systems.
Funding
This research is supported by the Natural Science Foundation of Chongqing (CSTB2024NSCQ-MSX0651) and the Team Building Project for Graduate Tutors in Chongqing (yds223010).
Data availability
The authors confirm that all data generated or analysed during this study are included in this article.
Declaration
Conflict of interest No potential conflict of interest was reported by the authors.